# What Matters in Hierarchical Search for Combinatorial Reasoning Problems?

### v1

@@ Appendix B Key factors for Hierarchical Search
  • presence of dead ends,
  • learning from diverse experts,
@@
  • complex action space
@@
B.1 Dead ends
In this paper, we conceptualize the environment $\mathcal{E}$ as a Markov Decision Process (MDP), where the environment, for any given state $s\in\mathcal{S}$ and action $a\in\mathcal{A}$, transitions to a subsequent state $s^{\prime}\in\mathcal{S}$. Within this framework, we define a dead-end state $s_{d}$ as a specific type of state with two defining characteristics:
  1. 1.
    The state $s_{d}$ does not represent a solution state.
  2. 2.
    From state $s_{d}$, there exists no feasible sequence of actions leading to a trajectory $(s_{d},s_{d+1},s_{d+2},\dots,s_{d+n})$ that culminates in a solution state $s_{d+n}$.
Dead-end states pose significant challenges in reinforcement learning, as they represent points from which no further learning or progress toward a solution can be made. Identifying and addressing these states is crucial for improving the efficiency and effectiveness of learning algorithms in navigating complex state spaces. Figure 14 shows an illustrative example of a dead-end state.
Examples of dead-ends in kSubS vs. BestFS
B.2 Learning from diverse experts
Usually broad datasets of offline demonstrations contain data from various sources [10, 26, 28]. Because of that, the demonstrations may be diverse in terms of length, structure, quality, skills, strategy, etc. We want to answer the question whether such setting is handled better by high-level or low-level search algorithms.
Experiment setup For this experiment, we choose the Rubik’s cube environment. We collected a dataset of 500 000500000500\,000500 000 trajectories, computed with four different solvers for the Rubik’s cube:
@@ B.2 Learning from diverse experts
Results As shown in Figure 2, the subgoal methods outperform the low-level methods by a wide margin. While BestFS is comparable on small budgets, it struggles with solving most of the instances. Also, it should be noted that the performance of the subgoal methods changes only slightly compared to training on a single Random solver (Figure 2) while the low-level searches are heavily affected.
@@
At the same time we observe that since BestFS and AdaSubS show nearly identical performance on Beginner solutions, it is questionable that hierarchical methods handle long-horizon tasks better, which is a common belief [43, 18, 7].
@@
B.3 Noisy value function
Accurate learning of value function presents a significant challenge in offline reinforcement learning, where the agent’s training is conducted on a fixed dataset of pre-collected trajectories. Section B.2 hints that when value estimates are subject to high uncertainty, subgoal methods should outperform low-level searches.
To confirm that intuition, we run an experiment in a Rubik’s cube environment (Section 4.1.4). This time, the training setup is the same as in Section 4.1.1. However, during inference, we add additional noise to the value estimates. That is, whenever a node is added to the search tree and its value estimate equals $\hat{v}$, we add it with the value of $\hat{v}+\mathcal{N}(0,\sigma)$ instead.
@@ B.3 Noisy value function
On the other hand, subgoal methods are much more resilient to noise in the value. Adaptive Subgoal Search is nearly not affected by the presence of noise. For subgoal distances longer than 1, kSubS retains as much as $40\%$ success rate, even with completely random values. The kSubS-1 variant behaves similarly to BestFS, for they are, in fact, equivalent.
@@
B.4 Complex action spaces
The primary difference between low-level methods and subgoal methods is that the former predicts the next action, and the latter – the next state. In many environments, the action space is as simple as a few bits, allowing for iterating over all possible actions, and sampling them. At the same time, states may be considerably larger, up to the extreme of image observations. However, in some environments, the action space is comparable to the state space, or even more complex. A classic example is the AntMaze environment, in which actions are 8-dimensional, while the goal space is only 2-dimensional [22].
Among the combinatorial reasoning environments we consider, INT has the most complex action space. In INT, actions correspond to proof steps and are represented as the chosen axiom, specification of its input entities, and the required premises [60]. Thus, the complexity of the action is at least comparable to the states. Moreover, solving the INT inequalities is based on constant simplification of the given expression, so the state is getting even smaller with each step.
Our experiments, shown in Figure 6, clearly confirm the advantage of using subgoal methods in the INT environment. To further verify the source of that advantage, we conducted another experiment, in a modified Rubik’s cube environment. Recall that the experiment presented in Section 4.1.1 shows that subgoals offer no significant advantage in the original Rubik’s cube. Now, we want to check whether the outcome would be different if the action space were more complex. For that purpose, we extended the action space 100 times. That is, the new action space consists of 1200 possible moves to choose from – 100 copies of each original action.
@@ Appendix C Network architectures & training details
We used BART [40] and BERT [16] architectures from HuggingFace Transformers for all components. Subgoal generators and INT’s policies (CLLP and baseline policy) use BART. The remaining policies and value functions use BERT. Following the practice in [63], we’ve reduced model size parameters, as detailed in Table 2.
INT As states in INT are complex objects, we prefer to use their string representations and avoid mapping arbitrarily generated strings into complex states. Requisite modifications to the component definition are best illustrated analogously to Section D.1. A generator is redefined as follows:
and conditional level policy:
Sokoban Unlike prior work [63, 14], which used convolutional networks for all components, we work on tokenized representations of Sokoban boards and use BERT/BART architectures instead. This modification did not adversely impact our ability to replicate AdaSubS and kSubS results.
Training pipeline We trained our models from scratch using the HuggingFace Transformer pipeline. Detailed training parameters, which varied across environments, can be found in Table 1.
@@ E.2 INT
Trajectories are constructed from sequences of axiom applications, similarly to [63], who followed [60]. A set of up to 15 (out of 18) axioms is first selected, and then a random axiom order is set and validated. Finally, a proof is converted to a relevant trajectory. Approximately 500,000 trajectories were generated for model pre-training.
@@ F.1 Best-First Search
In our implementation, we adhere to the Best-First Search principle by utilizing the learned value function, a common practice in the planning domain [6, 14, 63, 35]. It should be noted that in each of our experiments, all the compared algorithms use the same value function network. This way we ensure that the differences come solely from the algorithmic part.
@@
We note that the high-level algorithms could be made complete, as proposed in [36, 63]. However, to maximize the efficiency we choose to keep the tested algorithms in their original form. The ability to search with sparse trees not only lets the methods advance fast, but also withdraw quickly if the branch does not lead to the solution (is a dead end).
@@ F.2 Monte Carlo Tree Search
Our Monte Carlo Tree Search (MCTS) solver, designed for a single-player setting, is based on the AlphaZero framework [51]. The high-level workflow of MCTS is illustrated in Figure 29, and detailed pseudocode is provided in Algorithm 3.
@@
  • Simulation: Following the AlphaZero approach [51], policy and value networks replace traditional simulations. The policy network suggests favorable moves, while the value network predicts their probability of success, directing the algorithm towards beneficial trajectories.
@@ F.3 A Search
To tune the depth parameter for our experiments, we grided over values $[0.1,0.2,0.5,1,2,5,10]$. However, usually the best choice was to keep the cost weight low ($0.1$ or $0.2$, see Figure 30). While conservative search allows A* avoid more dead-ends than BestFS (see Table 8), usually greedy steps lead to finding the solution much faster.
Algorithm 5 Complete pseudocode for A*
Initial state $initial\_state$
Heuristic function $\mathcal{V}(node)$
function AStarSearch(($initial\_state$))
$OpenSet\leftarrow\{initial\_state\}$
$CameFrom\leftarrow$ an empty map
$GCost\leftarrow$ map with default value of $\infty$
$GCost[initial\_state]\leftarrow 0$
$FCost\leftarrow$ map with default value of $\infty$
$FCost[initial\_state]\leftarrow\mathcal{V}(initial\_state)$
while $OpenSet$ is not empty do
   $current\leftarrow$ node in $OpenSet$ with the lowest $FCost$ value
   if $current=goal\_state$ then
     return $ReconstructPath(CameFrom,current)$
@@
   $OpenSet.remove(current)$
   for each neighbor of $current$ do
     $TentativeGCost\leftarrow GCost[current]+dist\_between(current,neighbor)$
     if $TentativeGCost<GCost[neighbor]$ then
        $CameFrom[neighbor]\leftarrow current$
        $GCost[neighbor]\leftarrow TentativeGCost$
        $FCost[neighbor]\leftarrow GCost[neighbor]+\mathcal{V}(neighbor)$
        if $neighbor$ not in $OpenSet$ then
          $OpenSet.add(neighbor)$
        end if
@@
return failure
F.4 kSubS
kSubS can be seen as a special case of Adaptive Subgoal Search, but with just one generator $\rho_{k}$. For details see Algorithm 6. Both methods are introduced and studied in [14, 63].
F.5 AdaSubS
@@
F.6 HIPS and HIPS-$\varepsilon$
@@
   Use Best-First Search strategy to select a state $s$.
@@ Appendix I Further discussion on HIPS results
HIPS and HIPS-$\varepsilon$ [35, 36] are recent hierarchical search algorithms proposing to generate subgoals with variational autoencoders. We attempted to use HIPS and HIPS-$\varepsilon$ in greedy and prior-informed variations, and for all HIPS methods, the cost of inference was prohibitively high.
@@
Our evaluation, illustrated in Figure 34, shows that HIPS uses 100x more low-level nodes in search than comparable subgoal search methods and baselines - despite relatively similar subgoal efficiency as calculated in relevant papers. These findings informed our decision not to evaluate HIPS in the rest of the paper.

### v3

@@ Appendix B Key Factors For Hierarchical Search
  • learning from diverse data sources,
@@
  • complex action space,
  • presence of dead ends
@@
B.1 Learning from diverse data sources
Achieving superhuman performance in complex tasks, as demonstrated by AlphaGo Silver et al. (2016), often involves large-scale datasets of demonstrations obtained from agents with varying skill levels and strategies. However, this diversity introduces challenges such as inconsistencies in demonstrations and variations in quality (Fu et al., 2020; Chen et al., 2021; Levine et al., 2020). Widely used datasets like D4RL (Fu et al., 2020), Open X-Embodiment (Collaboration et al., 2023), or Waymo Open Dataset (Sun et al., 2020) reflect this diversity, highlighting the need to address these challenges effectively. We want to answer the question whether such setting is handled better by high-level or low-level search algorithms.
Experiment setup For this analysis, we focus on the Rubik’s cube environment. We collected a dataset of 500 000500000500\,000500 000 trajectories, computed with four different solvers for the Rubik’s cube:
@@ B.1 Learning from diverse data sources
Results As shown in Figure 3, the subgoal methods outperform the low-level methods by a wide margin. While $\rho$-BestFS is comparable on small budgets, it struggles with solving most of the instances. Also, it should be noted that the performance of the subgoal methods changes only slightly compared to training on a single Random solver (Figure 6) while the low-level searches are heavily affected.
@@
At the same time we observe that since BestFS and AdaSubS show nearly identical performance on Beginner solutions, it is questionable that hierarchical methods handle long-horizon tasks better, which is a common belief (Nachum et al., 2018; Eysenbach et al., 2019; Chen et al., 2024).
@@
B.2 Value Approximation Errors
In many practical scenarios, value function estimates are based on either limited data samples or handcrafted heuristics (Campbell et al., 2002; Mnih et al., 2015; Walke et al., 2023). This often leads to high approximation errors. If search algorithms rely too heavily on these imperfect estimates, they can make poor decisions, especially in large and complex environments where accurate value estimates are even harder to obtain (Collaboration et al., 2023; Vinyals et al., 2019).
Section B.1 hints that when value estimates are subject to high uncertainty, subgoal methods should outperform low-level searches. To confirm that intuition, we run an experiment in a Rubik’s cube, N-Puzzle, and Sokoban environments (Section 5.2). During inference, we add additional noise to the value estimates. That is, whenever a node is added to the search tree and its value estimate equals $\hat{v}$, we add it with the value of $\hat{v}+\mathcal{N}(0,\sigma)$ instead.
@@ B.2 Value Approximation Errors
On the other hand, subgoal methods are much more resilient to noise in the value. Adaptive Subgoal Search is nearly not affected by the presence of noise. kSubS is able to retain as much as $40\%-90\%$ success rate, even with completely random values.
@@
B.3 Complex Action Spaces
In environments with large action spaces, search methods often struggle due to the exponential increase in the number of choices at each decision point (Sutton & Barto, 1998). This complexity makes it difficult to efficiently identify optimal actions, slowing down decision-making and exploration (Dulac-Arnold et al., 2015; Silver et al., 2016).
The primary difference between low-level methods and subgoal methods is that the former predicts the next action, and the latter – the next state. In many environments, the action space is as simple as a few bits, allowing for iterating over all possible actions, and sampling them. At the same time, states may be considerably larger, up to the extreme of image observations. However, in some environments, the action space is comparable to the state space, or even more complex. A classic example is the AntMaze environment, in which actions are 8-dimensional, while the goal space is only 2-dimensional (Fu et al., 2020).
Among the combinatorial reasoning environments we consider, INT has the most complex action space. In INT, actions correspond to proof steps and are represented as the chosen axiom, specification of its input entities, and the required premises (Wu et al., 2021). Thus, the complexity of the action is at least comparable to the states. Moreover, solving the INT inequalities is based on constant simplification of the given expression, so the state is getting even smaller with each step.
Our experiments, shown in Figure 11, clearly confirm the advantage of using subgoal methods in the INT environment. To further verify the source of that advantage, we conducted another experiment, in a modified Rubik’s cube environment. Recall that the experiment presented in Section 5.1 shows that subgoals offer no significant advantage in the original Rubik’s cube (with a single data source). Now, we want to check whether the outcome would be different if the action space were more complex. For that purpose, we extended the action space 100 times. That is, the new action space consists of 1200 possible moves to choose from – 100 copies of each original action.
@@ B.3 Complex Action Spaces
B.4 Dead Ends
Dead-end states present a major challenge in decision-making and planning tasks. Once an agent encounters a dead end, reaching the goal becomes impossible, leading to wasted computational effort as the algorithm may continue exploring parts of the search space that do not contribute to solving the problem (Russell & Norvig, 2020). Failing to identify dead-ends may even lead to unsafe behavior (Fatemi et al., 2021; Sutton & Barto, 1998). At the same time, identifying dead-ends is NP-complete in many environments.
Specifically, a dead-end state $s$ is one from which there exists no feasible sequence of actions that leads to the goal state. Figure 31 shows an illustrative example of a dead-end state.
B.4.1 Examples Of Dead-Ends In kSubS vs. BestFS
@@ Appendix C Network Architectures & Training Details
We used BART (Lewis et al., 2020) and BERT (Devlin et al., 2019) architectures from HuggingFace Transformers for all components. Subgoal generators and INT’s policies (CLLP and baseline policy) use BART. The remaining policies and value functions use BERT. Following the practice in (Zawalski et al., 2023), we’ve reduced model size parameters, as detailed in Table 2.
INT As states in INT are complex objects, we prefer to use their string representations and avoid mapping arbitrarily generated strings into complex states. Requisite modifications to the component definition are best illustrated analogously to D.1. A generator is redefined as follows:
and conditional level policy:
Sokoban Unlike prior work (Zawalski et al., 2023; Czechowski et al., 2021), which used convolutional networks for all components, we work on tokenized representations of Sokoban boards and use BERT/BART architectures instead. This modification did not adversely impact our ability to replicate AdaSubS and kSubS results.
Training pipeline We trained our models from scratch using the HuggingFace Transformer pipeline. Detailed training parameters, which varied across environments, can be found in 1.
@@ E.2 INT
Trajectories are constructed from sequences of axiom applications, similarly to (Zawalski et al., 2023), who followed (Wu et al., 2021). A set of up to 15 (out of 18) axioms is first selected, and then a random axiom order is set and validated. Finally, a proof is converted to a relevant trajectory. Approximately 500,000 trajectories were generated for model pre-training.
@@ F.1 Best-First Search
In our implementation, we adhere to the Best-First Search principle by utilizing the learned value function, a common practice in the planning domain (Brunetto & Trunda, 2017; Czechowski et al., 2021; Zawalski et al., 2023; Kujanpää et al., 2023a). It should be noted that in each of our experiments, all the compared algorithms use the same value function network. This way we ensure that the differences come solely from the algorithmic part.
@@
We note that the high-level algorithms could be made complete, as proposed in (Kujanpää et al., 2023b; Zawalski et al., 2023). However, to maximize the efficiency we choose to keep the tested algorithms in their original form. The ability to search with sparse trees not only lets the methods advance fast, but also withdraw quickly if the branch does not lead to the solution (is a dead end).
@@ F.2 Monte Carlo Tree Search
Our Monte Carlo Tree Search (MCTS) solver, designed for a single-player setting, is based on the AlphaZero framework (Silver et al., 2018). The high-level workflow of MCTS is illustrated in Figure 35, and detailed pseudocode is provided in Algorithm 3.
@@
  • Simulation: Following the AlphaZero approach (Silver et al., 2018), policy and value networks replace traditional simulations. The policy network suggests favorable moves, while the value network predicts their probability of success, directing the algorithm towards beneficial trajectories.
@@ F.3 A Search
To tune the depth parameter for our experiments, we grided over values $[0.1,0.2,0.5,1,2,5,10]$. However, usually the best choice was to keep the cost weight low ($0.1$ or $0.2$, see Figure 36). While conservative search allows A* avoid more dead-ends than BestFS (see Figure 12), usually greedy steps lead to finding the solution much faster.
Algorithm 5 Complete pseudocode for $A^{*}$ Search
value function network $V$
policy $\rho_{BFS}$
predicate of solution solved
depth scaling factor $\lambda$
function search($s_{0}$)
$T\leftarrow\emptyset$priority queue}
$T$.push$((V(s_{0}),s_{0}))$
[…]
seen.add$(s_{0})$seen is a set}
[…]
   $\textit{actions}\leftarrow\rho_{BFS}(s)$
   for $a\textbf{ in }\textit{actions}$ do
     $s^{\prime}\leftarrow\textsc{envStep}(s,a)$
     if s $s^{\prime}\textbf{ in }\textit{seen}$ then
        continue
@@
     $\textit{seen}.\textsc{add}(s^{\prime})$
     $\textit{parents}[s^{\prime}]\leftarrow s$
     $T.\textsc{push}((V(s^{\prime})-\lambda\cdot\textit{depth$(s^{\prime})$},s^{\prime}))$
     if $(s^{\prime})$ then
        {solution found}
        return extractLowLevelTrajectory($s^{\prime}$, parents)
@@
return $\mathtt{False}$solution not found}
F.4 kSubS And AdaSubS
Overview
AdaSubS is a hierarchical search algorithm designed to solve combinatorial problems by operating on high-level nodes, which represent multiple steps rather than single actions. It employs multiple generators $\mathcal{G}_{k_{1}},\mathcal{G}_{k_{2}},\dots,\mathcal{G}_{k_{m}}$ to generate subsequent subgoals, a value function $\mathcal{V}$ to estimate the distance from a given state to the solution, and a conditional low-level policy $\mathcal{P}$ to execute a series of actions leading from one subgoal to the next. kSubS is a special case of AdaSubS, where only a single generator is used. These methods are introduced and studied in (Czechowski et al., 2021; Zawalski et al., 2023).
Stages
The method begins by adding $m$ initial nodes (one per each generator) to a priority queue, where each initial node $i$ is assigned a priority $(k_{i},\mathcal{V}(s_{0}))$. Here, $k_{i}$ is the length of the generator used during the node’s expansion, and $\mathcal{V}(s_{0})$ estimates the distance (in low-level actions) between $s_{0}$ and the solution. The following steps are repeated until a solution is found or the budget is exhausted:
  • Selection for expansion: The node $\left((k,\mathcal{V}(s),s\right)$ with the highest priority is extracted from the queue. This priority structure ensures that the algorithm prioritizes expanding the longest subgoals whenever possible.
  • Generating subgoals: The current state $s$ is passed to the selected generator $\mathcal{G}_{k}$, which produces multiple subgoal propositions represented as states $s^{*}_{1},s^{*}_{2},\dots,s^{*}_{p}$.
  • Verifying reachability: Since $\mathcal{G}_{k}$ can produce invalid or unreachable subgoals, each proposed subgoal must be verified. The conditional low-level policy $\mathcal{P}$ begins an iterative process, taking single steps from $s$ towards the proposed subgoal $s^{*}_{j}$. If $s^{*}_{j}$ is reached within $k$ steps, the subgoal is accepted, and new high-level nodes $\{\left((k_{i},\mathcal{V}(s^{*}_{j})),s^{*}_{j}\right)\}_{i\in\{1\dots m\}}$ are added to the priority queue as potential future subgoals to expand.
For a graphical overview of how AdaSubS works, see Appendix H.
@@
F.5 HIPS And HIPS-$\varepsilon$
@@
   Use PHS* search strategy to select a state $s$.
@@ Appendix I Further Discussion On HIPS Results
HIPS and HIPS-$\varepsilon$ (Kujanpää et al., 2023a; b) are recent hierarchical search algorithms proposing to generate subgoals with variational autoencoders. We attempted to use HIPS and HIPS-$\varepsilon$ in greedy and prior-informed variations, and for all HIPS methods, the cost of inference was prohibitively high.
@@
Our evaluation, illustrated in 40, shows that HIPS uses 100x more low-level nodes in search than comparable subgoal search methods and baselines - despite relatively similar subgoal efficiency as calculated in relevant papers. These findings informed our decision not to evaluate HIPS in the rest of the paper.